Grading Your grade in this course will be based on your performance on quizzes, three (3) mid-term exams, and a final exam. The weights assigned to each of these components will be:

Quizzes  10%
Mid-term exams (3) 60%
Final exam  30%

Course grades will be determined according to the following scale:

A 90 –100
B 80 – 89
C 70 – 79
D 60 – 69
F 0 – 59

Note that the deadline to drop this course with a grade of W is \underline{Friday, February 20, 1998}.

Exams There will be three (3) exams during the semester. Tentative dates and topics for these exams are:

Tuesday, February 10  Chapters 1 and 2
Thursday, March 5  Chapters 3 and 4
Thursday, April 16  Chapters 5 and 6

There will be no make-up exams. If you miss one exam due to a documented reason of illness, family emergency or participation in a University sponsored event, your score on the final exam will be used to replace the missing exam score. Excuses such as oversleeping, forgetting the time or location of the exam, and lack of studying are explicitly noted as unacceptable grounds for missing an exam.

A comprehensive final will be given at \underline{9:00a.m.} on \underline{Tuesday, May 5, 1998}.

Quizzes On average, there will be one quiz each week. Quiz questions will be based on the homework problems. Each quiz will be graded on a five (5) point scale. Your quiz grade will be based on the sum of your nine (9) highest quiz scores. There are no makeup quizzes.

Homework A minimum set of homework problems will be announced for each section that we discuss. We will discuss selected assigned problems in class, but no homework will be collected. You are responsible for working and understanding all of the assigned problems.

Attendance Regular class attendance is important. Consistent with the USC Undergraduate Bulletin, a grade penalty may be applied to any student missing more than three classes (10%) during the semester.

Academic Honesty Cheating and plagiarism will not be tolerated in this course. You will be working in groups in class and are encouraged to discuss homework problems with others. You will have to take all quizzes and exams on your own. Violations of this policy will be dealt with in a manner consistent with University guidelines.